OpenStudy (eric_d):

Express Partial fraction x^2+x-1/(x+1)(x+2)

OpenStudy (eric_d):

@Hero

OpenStudy (rational):

familiar with long division ?

OpenStudy (eric_d):

yes

OpenStudy (rational):

p(x) / q(x) degree of p(x) needs to be less than q(x) for partial fractions to work

OpenStudy (rational):

but in your rational function, both degrees are equal divide so that the degree of numerator becomes less than the denominator

OpenStudy (eric_d):

p(x) is the numerator

OpenStudy (eric_d):

I know this method

OpenStudy (luigi0210):

Oh, oh! Is this partial fraction decomposition? :D

OpenStudy (eric_d):

Yes

OpenStudy (luigi0210):

Okay, what you need to do is just completely ignore the top for now, and simplify the bottom as much as possible. So it's \(\Large \frac{x^2+x-1}{(x+1)(x+2)}\) right?

OpenStudy (eric_d):

ys

OpenStudy (luigi0210):

Alright, so now just break \(\Large \frac{x^2+x-1}{(x+1)(x+2)}\) into fractions with A and B and C's, etc, just as many as you need to match the denominator with a term: \[\Large \frac{x^2+x-1}{(x+1)(x+2)}=\frac{A}{(x+1)}+\frac{B}{(x+2)}\] ^That's how you would set it up as. Just a rule, whatever degree power you have at the bottom, the top will be one less of that. SO Since the bottom had a degree of 1 for x, the top would of been a power of 0, which is just 1, so A*1=A.

OpenStudy (luigi0210):

If you wanna solve it, just multiply it out by the common denominators

OpenStudy (eric_d):

...=A(x+2)+B(x+1)

OpenStudy (luigi0210):

Exactly! \(\Large x^2+x-1=A(x+2)+B(x+1)\) Use the left equation as a guide for values if you're gonna distribute out the right :)

OpenStudy (eric_d):

when x= -2 (-2)^2+(-2)-1=A(0)-B B=-1 when x=-1 A=1

OpenStudy (eric_d):

So,

OpenStudy (eric_d):

1/(x+1)- 1/(x+2)

OpenStudy (luigi0210):

If you plug in -1 you'll get \(\Large (-1)^2-1-1=A\), \(\Large (1-1)-1=A\) \(\Large 0-1=A\) So -1 for A too :P

OpenStudy (eric_d):

okay

OpenStudy (eric_d):

Hw wld the final answer be

OpenStudy (luigi0210):

What you just wrote with a -1 as A, or you could factor out that negative altogether \(\Large -(\frac{1}{x+1}+\frac{1}{x+2})\)

OpenStudy (eric_d):

Based on the answer from the book, it's 1- 1/(x+1)- 1/(x+2) So, is the book answer wrong ?

OpenStudy (eric_d):

?

OpenStudy (luigi0210):

No, the book isn't wrong, we just forgot to simplify in the beginning.. notice that if we FOIL, they will both have the same x power.

OpenStudy (luigi0210):

So we actually needed to composite this: \[\Large \frac{x^2}{x^2+3x+3}+\frac{x-1}{(x^2+3x+3)}\] Which is \(\Large 1+(\frac{x-1}{x^2+3x+3}) \)

OpenStudy (luigi0210):

So we just needed the simplification, in the beginning, everything else is good :)

OpenStudy (eric_d):

Okay, understood! Thanks Luigi

OpenStudy (luigi0210):

You're welcome, good luck bud~

hartnn (hartnn):

\(\Large \frac{x^2+x-1}{(x+1)(x+2)}=\frac{A}{(x+1)}+\frac{B}{(x+2)}\) that is wrong^^ as rational said, partial fractions do not work when numerator degree is EQUAL or HIGHER than denominator's

OpenStudy (luigi0210):

Yea I noticed that, we forgot the simplification in the beginning :l